### v2

# An accurate measurement of the spectral resolution of the JWST Near Infrared Spectrograph

An accurate measurement of the spectral resolution of the JWST Near Infrared Spectrograph
The spectral resolution () of spectroscopic data is crucial information for accurate kinematic measurements. In this letter, we present a robust measurement of the spectral resolution of the JWST’s Near Infrared Spectrograph (NIRSpec) in fixed slit (FS) and integral field spectroscopy (IFS) modes. Due to the similarity of the utilized slit dimension if the FS mode to that of the shutters in the multi-object spectroscopy (MOS) mode, our resolution measurements in the FS mode can also be used for the MOS mode in principle. We modeled H and He lines of the planetary nebula SMP LMC 58 using a Gaussian line spread function (LSF) to estimate the wavelength-dependent resolution for multiple disperser and filter combinations. We corrected for the intrinsic width of the planetary nebula’s H and He lines due to its expansion velocity by measuring it from a higher-resolution X-shooter spectrum. We find that NIRSpec’s in-flight spectral resolutions exceed the pre-launch estimates provided in the JWST User Documentation by 11–53% in the FS mode and by 1–24% in the IFS mode across the covered wavelengths. We recover the expected trend that the resolution increases with the wavelength within a configuration. The robust and accurate LSFs presented in this letter will enable high-accuracy kinematic measurements using NIRSpec for applications in cosmology and galaxy evolution.
Key Words.:
Methods: observational – Techniques: spectroscopic – Methods: data analysis1 Introduction
The JWST’s Near Infrared Spectrograph (NIRSpec; Böker et al., 2023) has enabled IR spectroscopy with an unprecedented combination of redshift reach, signal-to-noise ratio, and spatial resolution. One of its key applications is the study of stellar or gas kinematics in distant galaxies (e.g., de Graaff et al., 2024; Xu et al., 2024; Newman et al., 2025), helping us to understand their formation and evolution, the properties of gas outflows, and the impact of black hole and baryonic feedback (D’Eugenio et al., 2024). Furthermore, accurate measurement of the stellar velocity dispersion of lensing galaxies also enables high-precision measurement of the Hubble constant and other cosmological parameters through the method of time-delay cosmography (Shajib et al., 2025; TDCOSMO Collaboration, 2025).
Accurate measurement of the line spread function (LSF) or the spectral resolution , where is the full width at half maximum (FWHM), is a prerequisite to robustly measuring the velocity dispersion . The LSF induces additional broadening in the observed lines, which can be parametrized with assuming a Gaussian LSF. The accuracy requirement on to achieve a desired level of accuracy on can be estimated using the relation (Knabel et al., 2025)
Therefore, achieving a 1% accuracy in the velocity dispersion, as desirable for precision cosmology (Knabel et al., 2025), requires an estimate of NIRSpec’s medium-resolution accurate to 1.3% and 5.5% for km s-1 and 300 km s-1, respectively, for example.
NIRSpec’s LSF in the fixed-slit (FS) mode was previously measured from the narrow emission lines of the planetary nebula (PN) SMP LMC 58 (Isobe et al., 2023). These authors find that the resolution is higher by 10–20% than the pre-launch estimates provided by the JWST User Documentation (JDox).111JDox webpage for NIRspec dispersers and filters These authors assumed that the intrinsic expansion velocity of the PN is negligible. However, if the true expansion velocity is near the upper limit of typical values (i.e, 30 km s-1; Jacob et al., 2013), its impact may not be non-negligible, especially for the high-resolution grating (), which would require a 19% correction (2.7% for medium resolution with ). Nidever et al. (2024) find the resolution in the multi-object spectroscopy (MOS) mode’s G140H/F100LP configuration to be 50–85% higher than the JDox estimates, using narrow absorption lines in red giant stars. Through modeling with simulated data in the MOS mode, de Graaff et al. (2024) attribute the discrepancy between direct measurements and JDox estimates to differences in the source morphology. These authors demonstrate that a point source would lead to 50-90% higher resolution in the MOS G395H/F290LP configuration than the JDox estimates, which are based on uniformly illuminated slits. However, the predicted resolution by de Graaff et al. (2024) did not account for the additional broadening introduced by the data reduction procedure. Hence, this value is not directly applicable to the reduced data, and a direct measurement from real data remains indispensable.
In this letter, we provide a wavelength-dependent parametrization of NIRSpec’s resolution for multiple disperser–filter combinations for all of its FS, integral field spectroscopy (IFS), and MOS modes. We measured the resolutions by fitting prominent H and He emission lines for the PN SMP LMC 58 using datasets from JWST calibration programs, which were specifically obtained for wavelength and LSF calibrations. Isobe et al. (2023) also used a subset of this dataset for their LSF measurements. The same PN was also targeted by Jones et al. (2023) to measure the LSF for JWST’s Mid-Infrared Instrument (MIRI). Most of the fitted H lines have nearby weaker He multiplets, which are partially or fully blended and need to be accounted for to obtain an accurate characterization of the LSF. As a significant improvement over previous studies, we account for all the blended lines. We simultaneously fit the medium- and high-resolution spectra, thereby allowing lines that are deblended in the high-resolution spectrum to inform their relative strengths when fitting the medium-resolution one. Furthermore, we corrected for the PN’s expansion velocity after directly measuring it from a higher-resolution () spectrum of the same PN obtained using the Very Large Telescope’s (VLT) X-shooter instrument (Vernet et al., 2011).
This letter is organized as follows. In Section 2, we describe the JWST and X-shooter datasets we used. We describe our measurement method and provide the mesured values in Section 3, and then conclude the letter in Section 4. Code scripts and notebooks used in this analysis are publicly available on GitHub.222 https://github.com/ajshajib/nirspec_resolution.
2 Datasets
We describe the NIRSpec dataset used to measure the spectral resolutions in Section 2.1 and the VLT X-Shooter spectra used to measure the PN’s expansion velocity in Section 2.2.
2.1 NIRSpec spectra
To provide instrument calibrations for the NIRSpec, the planetary nebula SMP LMC 58 (also known as IRAS 052487007) was observed with the calibration program CAL-1492 (PI: T. Beck). We utilized the FS and IFS mode spectra from this program.333We do not utilize the MOS-mode observations from the program CAL-1492, as the exposures were stepped at 25 mas steps across the full pitch of an MSA shutter. Thus, the LSF in the stacked data is expected to have broadened beyond that from a point source centered on the slit with only sub-pixel dithers applied. However, due to the similarity of the S200A1 slit dimension to the MSA shutters, the FS-mode values are theoretically expected to be a reasonable proxy for MOS-mode observations. Suitable MOS-mode datasets in the future can be easily analyzed with the code suite and notebooks we provided. The FS mode used the S200A1 slit with 02 width, while including two primary dither positions with spectral sub-pixel dithers to improve the sampling. The IFS observations included a 4-point nod to improve sampling and enable background subtraction. We obtained the reduced Level-3 spectra from the Mikulski Archive for Space Telescopes (MAST) archive, which had been reduced with the default JWST pipeline v1.18.0 (Bushouse et al., 2023) with CRDS context jwst_1364.pmap. Thus, the reduced and stacked data followed the standard steps in the pipeline.
The FS-mode data were already extracted in the form of 1D spectra. For the IFS mode, we summed the spaxels within a circular aperture of radius 055, centered on the PN, to obtain the corresponding 1D spectra.
2.2 VLT X-shooter specrum
The VLT X-shooter spectrum of the PN was obtained as part of an Euclid-preparation program (program ID 110.23Q7.001; Euclid Collaboration et al., 2023). The spectrum was obtained with the VIS arm (covering 550–1020 nm) using a 12 slit, which has a nominal resolution of .444https://www.eso.org/sci/facilities/paranal/instruments/xshooter/inst.html We obtained the observed raw dataset from the European Southern Observatory (ESO) Science Portal,555https://archive.eso.org/scienceportal/ along with the calibration files. We reduced the dataset to produce an extracted 1D spectrum using ESO’s X-shooter data reduction pipeline v3.8.1 (Modigliani et al., 2010) with EsoReflex v2.11.5.
3 Measurement of the NIRSpec resolution
We adopted the Gaussian profile to characterize the NIRSpec LSF. By testing on a single emission line (Pa) that is not blended with other lines in both high and medium resolutions, we find that the LSF can be well-fitted with either a Gaussian or a Voigt profile, with the Voigt profile providing a better fit due to one additional degree of freedom. The Lorentzian profile, however, yields a significantly worse fit. We chose the Gaussian profile also because it is ubiquitously used to account for instrumental resolution in measurements of kinematics, for example, with the software program pPXF (Cappellari, 2017, 2023). The resolution parameter relates to the associated instrumental dispersion as , where is the speed of light. We parametrize the wavelength dependence of as
where is the instrumental dispersion not corrected for the PN’s expansion velocity , is the uncorrected instrumental dispersion at the pivot wavelength , and is the coefficient in the inverse-linear dependence on wavelength. We chose the inverse-linear -dependence for because the resolution is found approximately linear in from the pre-launch estimates, given that tends to be approximately constant in wavelength units. We describe our measurement process for in Section 3.1 and for in Section 3.2.
3.1 Emission line fitting from the NIRSpec spectra
We fit the spectra in wavelength slices around several prominent H and He emission lines. The details of the fitted wavelength ranges and the lists of lines within them are provided in Appendix A. Prominent H lines are often accompanied by He lines in proximity due to the resemblance in their atomic structures. Furthermore, He lines are usually multiplets. These lines are partially or fully blended, especially in the medium resolution. For a given combination of mode, disperser, and filter, we simultaneously fit all the lines within the chosen wavelength slices both in the high and medium resolution spectra, while individually constraining and for both gratings. The emission lines were fitted with post-pixelized (i.e., integrated within a pixel) Gaussian LSFs. The advantage of fitting all the lines together is that the information from all of them was aggregated to alleviate the issue that the LSF in most configurations is not critically sampled according to the Nyquist sampling theorem. The advantage of fitting both gratings together is that the relative line strengths could first be robustly constrained from the high-resolution spectrum, where they are less blended, which were then fixed for fitting the blended lines in the medium resolution (see Figure 1). We also fit for the PN’s relative motion (previously measured as km s-1; Reid & Parker, 2006) without any dependency on the wavelength. We allowed the PN velocity to be independent between the medium- and high-resolution spectra to account for any residual offset in the wavelength calibration. In each wavelength slice, we fit the continuum level with a linear function. We additionally allowed a Gaussian intrinsic scatter in the fitted to each line group around the mean relation of Eq. (2). Furthermore, we accounted for potential underestimation in the pipeline-produced uncertainties by allowing a systematic uncertainty floor to vary freely for each resolution. Thus, we have 10 non-linear parameters characterizing the global properties of our generative model: , , , , and for each resolution. The amplitudes of the line profiles and the continuum levels were solved through linear inversion. We obtained the posteriors of the non-linear parameters (provided in Table 1) using the nautilus sampler (Lange, 2023).
3.2 Measurement of the PN expansion velocity
To measure the expansion velocity of the PN from the X-shooter spectrum, we require a precise measurement of the spectral resolution of the X-shooter itself. We used the arc-lamp spectrum that was taken for wavelength calibration to obtain a wavelength-dependent estimate of X-shooter’s spectral resolution. We used the PypeIt software program (Prochaska et al., 2020) to measure the FWHMs of the arc-lamp lines in 15 wavelength bins within the covered wavelength range of 550–1020 nm (Figure 2). The uncertainty of the binned measurement was derived from the scatter in the individual-line measurements within each bin. The extent of the variation in the measured resolutions across the wavelength is qualitatively similar to that measured by Gonneau et al. (2020). We fit a quadratic function to the measured points. We also simultaneously fit a post-pixelized Gaussian function to the H line, which is the most prominent H line within the covered wavelengths. Although a few He lines also fall within the covered wavelength range, they do not significantly add to the constraining power due to their line strengths being smaller by 1.5 orders of magnitude or more. We parametrized the H line’s Gaussian profile with standard deviation , while allowing the peak position to vary to account for the PN’s relative motion. We constrain the PN’s expansion velocity km s-1.
We illustrate the measured resolution curves for all the mode–disperser combinations in Figure 3 and compare them to the corresponding pre-launch estimates provided by JDox. The resolutions in the IFS mode across the covered wavelengths are higher by 1–24% than the pre-launch estimates, whereas the resolutions in the FS mode are 11–53% higher.
| Disperser | Filter | () | (km s-1) | |
| FS mode | ||||
| G140M | F070LP | 10111.02 | ||
| G140H | F070LP | 10111.02 | ||
| G140M | F100LP | 13347.24 | ||
| G140H | F100LP | 13347.24 | ||
| G235M | F170LP | 22765.89 | ||
| G235H | F170LP | 22765.89 | ||
| G395M | F290LP | 38474.54 | ||
| G395H | F290LP | 38474.54 | ||
| IFS mode | ||||
| G140M | F100LP | 13539.59 | ||
| G140H | F100LP | 13539.59 | ||
| G235M | F170LP | 22938.52 | ||
| G235H | F170LP | 22938.52 | ||
| G395M | F290LP | 38891.71 | ||
| G395H | F290LP | 38891.71 |
4 Conclusion
In this letter, we measured NIRSpec’s spectral resolution in multiple configurations for in the FS and IFS modes. Due to the similarity in slit dimensions, the FS-mode resolution can be used as a reasonable proxy for the MOS mode. We provide a wavelength-dependent parametrization of the spectral resolution in terms of the instrumental dispersion by fitting multiple H and He emission lines across the observed wavelength range in the PN SMP LMC 58. We incorporated a correction for the intrinsic width of the PN’s lines, which we measured using a higher-resolution X-shooter spectrum of the same PN. Our measured resolutions are directly applicable for point source spectra; however, caution is advised when using them for extended sources.
The measured resolutions increase with wavelength in all configurations, as expected from the pre-launch estimates provided by JDox. For the FS mode, our measured resolution values extend up to 20% higher than those found by Isobe et al. (2023), potentially due to our methodology allowing multiple lines to be fitted within a blended line group. Our measured resolution in the G140H/F100LP configuration is smaller than that found by Nidever et al. (2024), which may be a difference between single-exposure and stacked spectra. Therefore, caution is advised when applying our measured value for datasets with drastically different dithering procedures from CAL-1492.
The parametric wavelength-dependent functions we provide will enable accurate measurements of kinematics from the NIRSpec data, as performed, e.g., by TDCOSMO Collaboration (2025) and Shajib et al. (2025). Although we do not provide the MOS-mode values at the medium resolution, the FS-mode values in medium resolution can be used as a substitute following Isobe et al. (2023), as the slit width for the analyzed data in the FS mode is the same as that of the multi-shutter-array slitlets.
Acknowledgements.
We thank David Nidever for helpful comments as the referee, which improved this manuscript. We thank Hsiao-Wen Chen, Danial Rangavar Langeroodi, Mark Morris, and Stefan Noll for useful discussions. This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with programs #1125 and 1492. The specific observations analyzed can be accessed via https://dx.doi.org/10.17909/m3c0-3t58. AJS and TT acknowledge support from NASA through STScI grants JWST-GO-2974, HST-GO-16773, and JWST-GO-1794. This research made use of numpy (Oliphant, 2015), scipy (Jones et al., 2001), astropy (Astropy Collaboration, 2013, 2018), jupyter (Kluyver et al., 2016), matplotlib (Hunter, 2007), seaborn (Waskom et al., 2014), and nautilus (Lange, 2023).References
- Astropy Collaboration (2013) Astropy Collaboration. 2013, A&A, 558, A33
- Astropy Collaboration (2018) Astropy Collaboration. 2018, AJ, 156, 123
- Böker et al. (2023) Böker, T., Beck, T. L., Birkmann, S. M., et al. 2023, Publications of the Astronomical Society of the Pacific, 135, 038001
- Bushouse et al. (2023) Bushouse, H., Eisenhamer, J., Dencheva, N., et al. 2023, JWST Calibration Pipeline
- Cappellari (2017) Cappellari, M. 2017, MNRAS, 466, 798
- Cappellari (2023) Cappellari, M. 2023, MNRAS, 526, 3273
- de Graaff et al. (2024) de Graaff, A., Rix, H.-W., Carniani, S., et al. 2024, A&A, 684, A87
- D’Eugenio et al. (2024) D’Eugenio, F., Pérez-González, P. G., Maiolino, R., et al. 2024, Nature Astronomy, 8, 1443
- Euclid Collaboration et al. (2023) Euclid Collaboration, Paterson, K., Schirmer, M., et al. 2023, A&A, 674, A172
- Gonneau et al. (2020) Gonneau, A., Lyubenova, M., Lançon, A., et al. 2020, A&A, 634, A133
- Hunter (2007) Hunter, J. D. 2007, Computing in Science and Engineering, 9, 90
- Isobe et al. (2023) Isobe, Y., Ouchi, M., Nakajima, K., et al. 2023, ApJ, 956, 139
- Jacob et al. (2013) Jacob, R., Schönberner, D., & Steffen, M. 2013, A&A, 558, A78
- Jones et al. (2001) Jones, E., Oliphant, T., Peterson, P., & Others. 2001, SciPy: Open Source Scientific Tools for Python
- Jones et al. (2023) Jones, O. C., Álvarez-Márquez, J., Sloan, G. C., et al. 2023, MNRAS, 523, 2519
- Kluyver et al. (2016) Kluyver, T., Ragan-Kelley, B., Pérez, F., et al. 2016, in Positioning and Power in Academic Publishing: Players, Agents and Agendas, ed. F. Loizides & B. Schmidt (IOS Press BV, Amsterdam, Netherlands), 87–90
- Knabel et al. (2025) Knabel, S., Mozumdar, P., Shajib, A. J., et al. 2025, TDCOSMO XIX: Measuring Stellar Velocity Dispersion with Sub-Percent Accuracy for Cosmography
- Lange (2023) Lange, J. U. 2023, MNRAS, 525, 3181
- Modigliani et al. (2010) Modigliani, A., Goldoni, P., Royer, F., et al. 2010, in Observatory Operations: Strategies, Processes, and Systems III, Vol. 7737, 773728
- Newman et al. (2025) Newman, A. B., Gu, M., Belli, S., et al. 2025, arXiv e-prints, arXiv:2503.17478
- Nidever et al. (2024) Nidever, D. L., Gilbert, K., Tollerud, E., et al. 2024, in IAU Symposium, Vol. 377, Early Disk-Galaxy Formation from JWST to the Milky Way, ed. F. Tabatabaei, B. Barbuy, & Y.-S. Ting, 115–122
- Oliphant (2015) Oliphant, T. E. 2015, Guide to NumPy, 2nd edn. (USA: CreateSpace Independent Publishing Platform)
- Prochaska et al. (2020) Prochaska, J., Hennawi, J., Westfall, K., et al. 2020, JOSS, 5, 2308
- Reid & Parker (2006) Reid, W. A. & Parker, Q. A. 2006, Monthly Notices of the Royal Astronomical Society, 373, 521
- Shajib et al. (2025) Shajib, A. J., Treu, T., Suyu, S. H., et al. 2025, arXiv e-prints, arXiv:2506.21665
- TDCOSMO Collaboration (2025) TDCOSMO Collaboration. 2025, arXiv e-prints, arXiv:2506.03023
- van Hoof (2018) van Hoof, P. A. M. 2018, Galaxies, 6, 63
- Vernet et al. (2011) Vernet, J., Dekker, H., D’Odorico, S., et al. 2011, A&A, 536, A105
- Waskom et al. (2014) Waskom, M., Botvinnik, O., Hobson, P., et al. 2014, Seaborn: V0.5.0 (November 2014)
- Xu et al. (2024) Xu, Y., Ouchi, M., Yajima, H., et al. 2024, ApJ, 976, 142
Appendix A Fitted line lists
In this appendix, we provide the list of fitted lines within the chosen wavelength ranges in Tables 2, 3, and 4 for the G140, G235, and G395 dispersers, respectively. We first obtained a comprehensive list of nebular H and He lines from the Atomic Line List database777v.300b5: https://linelist.pa.uky.edu/newpage/ (van Hoof 2018). We then aggregated He lines that are close to each other within 1/8th of a pixel size in the corresponding high-resolution spectrum into a single line with an average wavelength weighted by the transition probability (i.e., the Einstein coefficient ). We adopted 1/8 of the pixel size here, as this is half of the target threshold for wavelength calibration error (Böker et al. 2023). Thus, lines that are separated more closely can be sufficiently considered indistinguishable in the spectrum. Not all the line groups were fitted in some modes, as those line groups fall fully or partially outside of the covered wavelength range in the particular modes. The tables in this appendix specify the modes for which each line group was fitted.
| Principal Line | () | () | H lines () | He lines () | Mode | ||||
| Pa | 8980 | 9060 |
FS (070LP) | ||||
| Pa | 9190 | 9290 |
FS (070LP) | ||||
| Pa | 10020 | 10100 |
FS (F100LP), IFS | ||||
| He | 10780 | 10900 | – |
FS (F100LP), IFS | ||||
| Pa | 10910 | 10990 |
FS, IFS | ||||
| Pa | 12760 | 12885 |
FS (F100LP), IFS | ||||
| Br | 17350 | 17430 |
FS (F100LP), IFS | ||||
| Pa | 18675 | 18820 |
IFS |
| Principal Line | () | () | H lines () | He lines () | Mode | ||||
| Br | 16800 | 16850 |
IFS | ||||
| Br | 17300 | 17440 |
FS, IFS | ||||
| Pa | 18655 | 18840 |
FS, IFS | ||||
| He | 20555 | 20660 | – |
FS, IFS | ||||
| Br | 21575 | 21750 |
FS, IFS | ||||
| Br | 26180 | 26350 |
FS, IFS | ||||
| Pf | 30350 | 30480 |
IFS |
| Principal Line | () | () | H lines () | He lines () | Mode | ||||||||
| Pf | 30330 | 30520 |
FS, IFS | ||||||||
| Pf | 37350 | 37580 |
FS, IFS | ||||||||
| Br | 40300 | 40730 |
FS, IFS | ||||||||
| He | 42900 | 43100 | – |
FS, IFS | ||||||||
| Hu | 43725 | 43900 |
FS, IFS | ||||||||
| Hu | 51230 | 51450 |
IFS |